# Does GenAI Make Usability Testing Obsolete?

### v1

@@ VII Discussion and Conclusion
Our results show that—with fairly low effort and widely-available models—GenAI can predict in source code valid usability issues, that can easily be reviewed and fixed before releasing the app, avoiding to dissatisfy users and compromise their experience. This saves time and resources particularly for smaller app teams and freelancers, who tend to ignore usability evaluation [41]. However, our study also shows that relevant issues particularly identified by usability experts were missed by UX-LLM. This suggests that traditional usability evaluation can certainly be boosted with GenAI but not completely replaced, still requiring a UX expert in the loop.

### v2

@@ VII Discussion and Conclusion
Our results show that—with fairly low effort and widely-available models—GenAI can predict in source code valid usability issues, that can easily be reviewed and fixed before releasing the app, avoiding to dissatisfy users and compromise their experience. This saves time and resources particularly for smaller app teams and freelancers, who tend to ignore usability evaluation [40]. However, our study also shows that relevant issues particularly identified by usability experts were missed by UX-LLM. This suggests that traditional usability evaluation can certainly be boosted with GenAI but not completely replaced, still requiring a UX expert in the loop [70].
@@
Acknowledgement
This work was partly supported by MaibornWolff GmbH. We thank all participants in the usability testing, the UX experts, as well as the participants of the focus group for their time and feedback.